a farmer keeps hens and rabbits in his farm.one day he counted a total of 70 heads and 196 legs.how many more hens does he has then rabbits?
Answer by Menjax(62)

Let h = number of hens and r = number of rabbits
h+r=70 (total heads = total number of animals)(equation 1)
2h+4r=196 (each hen has 2 legs and rabbits have 4)(equation 2)
Times equation #1 by 2
2h+2r=140 (equation 3)
Equation 2 minus equation 3
2r=56
r=28
h=42
There are 14 more hens than there are rabbits